OpenStudy (anonymous):

u + 3v − 2w, u = <4, 2>, v = <-6, 1>, w = <2, 0>

OpenStudy (zzr0ck3r):

it might be easier to see these with different notation write your vectors as \(u=4i+2j\\v=-6i+1j\\w=2i\) then \(u+3v-2w=(4i+2j)+3(-6i+j)-2(2i)\) now collect like terms \(4i+2j-18i+3j-4i=-18i+5j=\langle -18,5\rangle\)